### v2

@@ 1 Introduction
Position-based cryptography (PBC), initially introduced by Chandran, Goyal, Moriarty and Ostrovsky [CGMO09], aims to allow a party to use its geographical location as a credential to implement various cryptographic protocols. An important building block for PBC is so-called Position Verification (PV), where an untrusted prover $P$ wants to convince a set of verifiers $V_{0},\dots,V_{k}$ that she is at a certain position $r$. In [CGMO09] it was proven that no secure classical protocol for position verification can exist, since there exists a general attack based on copying classical information. Adrian Kent first studied PV protocols that use quantum information in 2002, originally named quantum tagging [AKB06, KMS11] and currently called Quantum Position Verification (QPV) in the literature.Due to the no-cloning theorem [WZ82], attacks based on simply copying information do not transfer to the quantum setting. Nevertheless, unconditionally secure QPV was proven to be impossible, with [BCF+14] showing that any QPV could be attacked with double-exponential entanglement, which was later improved to exponential entanglement [BK11].
@@
On the other side, it is possible to prove a linear lower bound for the amount of entanglement [BK11, TFKW13] – an exponential gap. Still, the extreme inefficiency of the general attack left the possibility open that QPV could be shown secure if we consider attackers that are bounded in some (realistic) way. Ideally that would mean a protocol which requires an exponential amount of entanglement to attack, but any provable gap between the hardness of executing the protocol versus breaking the protocol could be interesting. This possibility created the opportunity for many interesting follow-up work, showing security under model assumptions such as finding clever polynomial attacks to proposed protocols [LL11, CL15, Spe16a, Dol19, DC22, GC19, CM22] and analyzing security in other models, such as the random oracle model [Unr14, LLQ22, GLW16].
@@
Particularly interesting here is the recent work by Liu, Liu, and Qian [LLQ22], which surprisingly shows that classical communication suffices to construct a secure protocol – immediately solving the issue of transmission loss. Despite this protocol being a great theoretical breakthrough, there are some downsides which make it unappealing to implement in the near future. A minor theoretical downside is that this proposed protocol is secure only under computational assumptions, and additionally requires the random oracle model to be secure against any entanglement. The larger practical downside is that the honest prover in [LLQ22]’s protocol requires a large quantum computer to execute the protocol’s steps – instead of manipulating and measuring single qubits as in most other protocols we’re considering.
@@
One of the simplest and best-studied QPV protocols, which constitutes the basis of this work, is based on BB84 states, which we will denote $\mathrm{QPV}_{\mathrm{BB84}}$. This protocol was initially introduced in [KMS11] and later on proved secure against unentangled adversaries in [BCF+14], with parallel repetition shown in [TFKW13] [^2] The analysis was tightened in [RG15], however this was proven in a slightly-weaker security model where the attackers share a round of classical communication instead of quantum communication, so the analyses are not directly comparable.. The protocol, explained in detail below, consists of one verifier ($V_{0}$) sending a BB84 state to the prover and the other verifier ($V_{1}$) sending a classical bit describing in which basis the prover has to measure, either the computational basis or the Hadamard basis. The prover then has to broadcast the measurement outcome to both verifiers, with all communication happening at the speed of light. This protocol can be attacked by attackers that share a single EPR pair. However, by encoding the basis information in several longer classical strings $x,y\in\{0,1\}^{n}$ that have to be combined in some way (i.e., the basis is given by $f(x,y)$ for some function $f$) it is possible to construct a single-qubit protocol, $\mathrm{QPV}_{\mathrm{BB84}}^{f}$, for which the quantum effort to break it grows with the amount of classical information that an honest party would use. This type of extension was already considered by Kent, Munro, and Spiller [KMS11] for a slightly different protocol, and analyzed more in-depth by Buhrman, Fehr, Schaffner, and Speelman [BFSS13], with a linear lower bound shown by Bluhm, Christandl, and Speelman [BCS22] [^3] This linear lower bound is not necessarily tight – the best currently-known attack on $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  requires exponential entanglement in the number of classical bits for most functions $f$..(Also see [JKPPG22] for a related lower bound for a slightly different protocol which combines quantum and classical communication.) The ability to enhance security by adding classical information makes the functional version of $\mathrm{QPV_{BB84}}$ an appealing candidate for future use.
@@
However, applying QPV experimentally encounters implementation problems, of which two are large enough that they force us to redesign our protocols. Whereas the transmission of classical information without loss at the speed of light is technologically feasible, e.g. via radio waves, the quantum counterpart faces obstacles. Firstly, most QPV protocols require the quantum information to be transmitted at the speed of light in vacuum, but the speed of light in optical fibers, a medium which would be necessary for many practical scenarios, is significantly lower than in vacuum. Secondly, a sizable fraction of photons is lost in transmission in practice. For this loss problem, we can distinguish two recent approaches. The first of which is to create protocols which are secure against any amount of loss, which we can call fully loss-tolerant protocols. This type of protocol was first introduced by Lim, Xu, Siopsis, Chitambar, Evans, and Qi [LXS+16], based on ideas from device-independent QKD. New examples and further analyses of fully loss-tolerant protocols were given by Allerstorfer, Buhrman, Speelman, and Verduyn Lunel [ABSV21, ABSVL22]. These protocols could be excellent realistic candidates for an implementation of QPV, but in the longer term they have two shortcomings: they are not secure against much entanglement – [ABSVL22] for instance show that if security against unbounded loss is required, this is unavoidable – and they require fast transmission of quantum information.
@@
In this paper, we study the security of the $\mathrm{QPV}_{\mathrm{BB84}}$  protocol in the lossy case, for attackers that do not pre-share entanglement, but that are allowed to perform local operations and a single simultaneous round of quantum communication (LOBQC) [GC19]. The best bounds for the non-lossy version of this protocol (for attackers that are allowed a round of quantum communication) come through reducing them to a monogamy-of-entanglement game [TFKW13] – we extend this game to include a third ‘loss’ response. Via Semidefinite Programming (SDP), we can then tightly bound the game’s winning probability, and therefore $\mathrm{QPV}_{\mathrm{BB84}}$’s security under photon loss, showing that a naive mixture of the extremal strategies turns out to give the optimal combination of error and response rate for the attackers [^4] Note that in our current work we solve the semidefinite programs for a complete range of experimental error probabilities, thereby obtaining an exhaustive characterization of the protocols we consider. For an experimental implementation only a small number of SDPs would have to be solved, namely the one that corresponds to a small interval around the error and loss present in the experimental set-up, necessitating much less computation time than used in preparing our results.. This is a three-player game involving two players, who have complete freedom, and a referee who performs a fixed measurement for every input. It is well-known how to analyze two-party scenarios using SDP relaxations, but extending these methods to three-party scenarios presents an obstacle. To obtain SDP bounds, we combine the NPA hierarchy [NPA08] with extra inequalities we derive from the relation between the referee’s measurements – the way these inequalities enable us to tackle a three-player problem using SDPs might be of independent interest.
Importantly, we are able to also show that these results can be adapted to show bounds for the lossy version of $\mathrm{QPV}_{\mathrm{BB84}}^{f}$. We do this by defining a new relaxation of our earlier SDP, which holds for the unentangled case, and show that the numerical bounds for this new SDP can be used to reprove a key lemma in [BCS22] for the case of our protocol. For instance, if the measurement error is very low, our results show that the single-qubit protocol $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  remains secure against attackers that share an amount of entanglement that is sublinear in the amount of classical bits, even when only, e.g., $51\%$ of the photons arrive at the honest party. The security proof even goes through in the case that the transmission of quantum information is slow, as long as classical information can be transmitted fast.
@@
The main result of this paper can be summarized as follows (see Theorems 3.12 and 5.7 for a formal version):
@@
(Informal). If some attackers pre-share a linear amount (in the size of the classical information) of qubits before the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$/ $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ protocol and if they respond with probability $\eta$, then the probability that they do not answer correctly is strictly larger than a numerical bound, which is a function of $\eta$. Therefore, if an honest prover, for which qubits arrive with probability $\eta$, has lower error probability, the protocol is correct and secure.
@@
Our work is directly related to the questions asked by Qi and Siopsis [QS15]. The most important difference between that work and ours, is that we are considering a more general security model. The attack analysis in [QS15] is based upon an assumption that attackers would immediately measure an incoming qubit using a projective measurement, and then communicate classically. By reducing to (extensions of) monogamy-of-entanglement games, we instead capture any quantum action of the attackers, including a round of quantum communication – it’s shown in [ABSVL22] that this stronger security model can in fact make a difference in the security. Critically, this also means that our method of analysis extends to the case where a classical function is combined with a single qubit, and our results make it possible to also prove entanglement bounds, instead of just bounds against unentangled protocols. Finally, our results can be reapplied in other settings that can be described by a monogamy-of-entanglement game.
There are two other recent papers by Allerstorfer, Buhrman, Speelman, and Verduyn Lunel [ABSV21, ABSVL22] that study the role of loss in quantum position verification. Those works provide many new results on the capabilities and limitations of fully loss-tolerant protocols.
Our work builds upon the (unpublished) results of Buhrman, Schaffner, Speelman, and Zbinden [Spe16b, Chapter 5]. That work took the initial step to bound QPV protocols in the lossy case using SDPs. However, its analysis was incomplete and numerically significantly less tight, e.g., its bounds on $\mathrm{QPV_{BB84}}$ were a factor two worse than the current work. Additionally, we significantly extend its applicability by extending the technique to also show entanglement bounds for the $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  protocol [BFSS13, BCS22].
2 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol and its security under unentangled attackers
@@
(A round of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol). Let $\eta$ be the transmission rate of the quantum channel, we define one round of the lossy-BB84 QPV protocol, denoted by $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, as follows (see Fig. 1 for steps 2. and 3.):
@@
    $V_{0}$ and $V_{1}$ secretly agree on a random bit $z$ and a basis $x\in\{0,1\}=:\mathcal{X}$, i.e. the computational $(0)$ or the Hadamard $(1)$ basis. In addition, $V_{0}$ prepares the EPR pair $\left|\Omega\right\rangle=(\left|00\right\rangle+\left|11\right\rangle)/\sqrt{2}$.
@@
    $V_{0}$ sends one qubit $Q$ of $\left|{\Omega}\right\rangle$ to $P$, and $V_{1}$ sends $x$ to $P$, coordinating their times so that $Q$ and $x$ arrive at $P$ at the same time.
@@
    Immediately, $P$ measures $Q$ in the basis $x$ and broadcasts her outcome, either 0 or 1, to $V_{0}$ and $V_{1}$. If the photon is lost, she sends $\perp$. Therefore, the possible answers from $P$ are $a\in\{0,1,\perp\}$.
@@
      both $a$ and $b$ arrive on time and $a=b$, and if
      • they equal $z$, the verifiers output ‘Correct’,
      • they equal $1-z$, the verifiers output ‘Wrong’,
      • they equal $\perp$, the verifiers output ‘No photon’,
@@
      if either $a$ or $b$ do not arrive on time or $a\neq b$, the verifiers output ‘Abort’.
The protocol is run sequentially $r$ times. Let $r_{\textsc{C}}$, $r_{\textsc{W}}$, $r_{\textsc{N}}$ and $r_{\textsc{A}}$ denote the number of times that the verifiers output ‘Correct’, ‘Wrong’, ‘No photon’ and ‘Abort’, respectively, after the $r$ rounds. The verifiers accept the prover’s location if the answers they receive match with $\eta$ and $p_{err}$ and they do not receive any different or delayed answers, i.e. if
$\begin{split}r_{\textsc{C}}&\approx r\eta(1-p_{err}),\hskip 42.67912ptr_{\textsc{W}}\approx r\eta(1-p_{err}),\\ r_{\textsc{N}}&\approx r(1-\eta),\hskip 62.59596ptr_{\textsc{A}}=0.\end{split}$ (1)
The symbols $\approx$ are due to the fact that in an implementation the above values would only be reached for $r\to\infty$. Notice that the verifiers remain strict with $r_{\textsc{A}}$, since an honest party would never send different answers and using current technology, her answers will arrive on time.
The $\mathrm{QPV}_{\mathrm{BB84}}$  protocol is recovered for $\eta=1$ and $p_{err}=0$. The above description of one round of the protocol corresponds to the purified version of the originally stated $\mathrm{QPV}_{\mathrm{BB84}}$  (for $\eta=1$). Notice that sending a single qubit $z$ in the basis $x$, as in the original version, instead of a qubit from an EPR pair, is completely equivalent to the above description. When $V_{0}$ measures $Q$ in the basis $x$, obtaining the random bit $z$, the reduced state sent to $P$ is given by $H^{x}\left|z\right\rangle$, where $H$ is the Hadamard matrix, and the original version is recovered. We will use the purified version in the analysis, because the security proofs are easier (in particular, it makes it easy to delay the choice of basis $x$ to a later point in time, which makes it clear that the attackers’ actions are independent of the basis choice).
@@
Figure 1: Steps 2. and 3. of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol, where straight lines represent classical information and undulated lines represent quantum information. Image reproduced from [BCS22] with permission.
@@
Whereas it is well-known that unlimited attackers can always successfully break any QPV protocol [BCF+14], the proof of the security of the $\mathrm{QPV_{BB84}}$ protocol under attackers that do not pre-share entanglement [BCF+14] opened a branch of study. The most general attack to a 1-dimensional QPV protocol is to place an adversary between $V_{0}$ and the prover, which we will call Alice, and another adversary between the prover and $V_{1}$, which we will call Bob. It is easy to see that having more than two adversaries in a 1-dimensional setting does not improve an attack. We now describe a general attack for attackers that do not pre-share entanglement.
Attack 2.2.
The most general attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, under the restriction that the attackers do not pre-share entanglement and are allowed to perform local operations and quantum communication (LOQC), consists of:
@@
    Alice intercepts the qubit $Q$ and applies an arbitrary quantum operation to it, and possibly some ancillary systems she possesses, ending up with the state $\rho$. She keeps a part of it and sends the other to Bob. On the other side, Bob intercepts $x$, copies it and sends the copy to Alice. Since they share no entanglement, any quantum operation that Bob could perform as a function of $x$ can be included in Alice’s operation (see e.g., [BCF+14, TFKW13]).

### v3

@@ 1 Introduction
Position-based cryptography (PBC), initially introduced by Chandran, Goyal, Moriarty and Ostrovsky [CGMO09], aims to allow a party to use its geographical location as a credential to implement various cryptographic protocols. An important building block for PBC is so-called Position Verification (PV), where an untrusted prover P wants to convince a set of verifiers $V_{0},\dots,V_{k}$ that she is at a certain position $r$. In [CGMO09] it was proven that no secure classical protocol for position verification can exist, since there exists a general attack based on copying classical information. Adrian Kent first studied PV protocols that use quantum information in 2002, originally named quantum tagging [KMSB06, KMS11] and currently called Quantum Position Verification (QPV) in the literature. Due to the no-cloning theorem [WZ82], attacks based on simply copying information do not transfer to the quantum setting. Nevertheless, unconditionally secure QPV was proven to be impossible, with [BCF+14] showing that any QPV could be attacked with double-exponential entanglement, which was later improved to exponential entanglement [BK11].
@@
On the other side, it is possible to prove a linear lower bound for the amount of entanglement [BK11, TFKW13]—an exponential gap. Still, the extreme inefficiency of the general attack left the possibility open that QPV could be shown secure if we consider attackers that are bounded in some (realistic) way. Ideally that would mean a protocol which requires an exponential amount of entanglement to attack, but any provable gap between the hardness of executing the protocol versus breaking the protocol could be interesting. This possibility created the opportunity for many interesting follow-up work, showing security under model assumptions such as finding clever polynomial attacks to proposed protocols [LL11, CL15, Spe16a, Dol22, DC22, GC20, CM23] and analyzing security in other models, such as the random oracle model [Unr14, LLQ22, GLW16].
@@
Particularly interesting here is the recent work by Liu, Liu, and Qian [LLQ22], which surprisingly shows that classical communication suffices to construct a secure protocol—immediately solving the issue of transmission loss. Despite this protocol being a great theoretical breakthrough, there are some downsides which make it unappealing to implement in the near future. A minor theoretical downside is that this proposed protocol is secure only under computational assumptions, and additionally requires the random oracle model to be secure against any entanglement. The larger practical downside is that the honest prover in [LLQ22]’s protocol requires a large quantum computer to execute the protocol’s steps—instead of manipulating and measuring single qubits as in most other protocols we’re considering.
@@
One of the simplest and best-studied QPV protocols, which constitutes the basis of this work, is based on BB84 states, which we will denote $\mathrm{QPV}_{\mathrm{BB84}}$. This protocol was initially introduced in [KMS11] and later on proved secure against unentangled adversaries in [BCF+14], with parallel repetition shown in [TFKW13] [^1] The analysis was tightened in [RG15], however this was proven in a slightly-weaker security model where the attackers share a round of classical communication instead of quantum communication, so the analyses are not directly comparable.. The protocol, explained in detail below, consists of one verifier ($V_{0}$) sending a BB84 state to the prover and the other verifier ($V_{1}$) sending a classical bit describing in which basis the prover has to measure, either the computational basis or the Hadamard basis. The prover then has to broadcast the measurement outcome to both verifiers, with all communication happening at the speed of light. This protocol can be attacked by attackers that share a single EPR pair. However, by encoding the basis information in several longer classical strings $x,y\in\{0,1\}^{n}$ that have to be combined in some way (i.e. the basis is given by $f(x,y)$ for some function $f$) it is possible to construct a single-qubit protocol, $\mathrm{QPV}_{\mathrm{BB84}}^{f}$, for which the quantum effort to break it grows with the amount of classical information that an honest party would use. This type of extension was already considered by Kent, Munro, and Spiller [KMS11] for a slightly different protocol, and analyzed more in-depth by Buhrman, Fehr, Schaffner, and Speelman [BFSS13], with a linear lower bound shown by Bluhm, Christandl, and Speelman [BCS22] [^2] This linear lower bound is not necessarily tight—the best currently-known attack on $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  requires exponential entanglement in the number of classical bits for most functions $f$..(Also see [JKPPG22] for a related lower bound for a slightly different protocol which combines quantum and classical communication.) The ability to enhance security by adding classical information makes the functional version of $\mathrm{QPV_{BB84}}$ an appealing candidate for future use.
@@
However, applying QPV experimentally encounters implementation problems, of which two are large enough that they force us to redesign our protocols. Whereas the transmission of classical information without loss at the speed of light is technologically feasible, e.g. via radio waves, the quantum counterpart faces obstacles. Firstly, most QPV protocols require the quantum information to be transmitted at the speed of light in vacuum, but the speed of light in optical fibers, a medium which would be necessary for many practical scenarios, is significantly lower than in vacuum. Secondly, a sizable fraction of photons is lost in transmission in practice. In practice, if telecom wavelength ($\sim$1550nm) single-photon sources are used, and the photons are sent through optical fibers with a loss of $0.15$dB/km [CZD19]. For this loss problem, we can distinguish two recent approaches. The first of which is to create protocols which are secure against any amount of loss, which we can call fully loss-tolerant protocols. This type of protocol was first introduced by Lim, Xu, Siopsis, Chitambar, Evans, and Qi [LXS+16], based on ideas from device-independent QKD. New examples and further analyses of fully loss-tolerant protocols were given by Allerstorfer, Buhrman, Speelman, and Verduyn Lunel [ABSL22, ABSVL22]. These protocols could be excellent realistic candidates for an implementation of QPV, but in the longer term they have two shortcomings: they are not secure against much entanglement—[ABSVL22] for instance show that if security against unbounded loss is required, this is unavoidable—and they require fast transmission of quantum information.
@@
In this paper, we study the security of the $\mathrm{QPV}_{\mathrm{BB84}}$  protocol in the lossy case in the No Pre-shared Entanglement (No-PE) model [BCF+14]—that is, for attackers that do not pre-share entanglement prior to the execution of the protocol, but that are allowed to perform local operations and a single simultaneous round of quantum communication (LOBQC) [GC20]. Security in this model is motivated by the fact that generating and distributing high-quality entanglement over long distances remains technologically challenging, so the No-PE model captures currently practical limitations. The best bounds for the non-lossy version of this protocol in the No-PE model come through reducing them to a monogamy-of-entanglement game [TFKW13]—we extend this game to include a third ‘loss’ response. Via Semidefinite Programming (SDP), we can then tightly bound the game’s winning probability, and therefore $\mathrm{QPV}_{\mathrm{BB84}}$’s security under photon loss, showing that a naive mixture of the extremal strategies turns out to give the optimal combination of error and response rate for the attackers [^3] Note that in our current work we solve the semidefinite programs for a complete range of experimental error probabilities, thereby obtaining an exhaustive characterization of the protocols we consider. For an experimental implementation only a small number of SDPs would have to be solved, namely the one that corresponds to a small interval around the error and loss present in the experimental set-up, necessitating much less computation time than used in preparing our results.. This is a three-player game involving two players, who have complete freedom, and a referee who performs a fixed measurement for every input. It is well-known how to analyze two-party scenarios using SDP relaxations, but extending these methods to three-party scenarios presents an obstacle. To obtain SDP bounds, we combine the NPA hierarchy [NPA08] with extra inequalities we derive from the relation between the referee’s measurements—the way these inequalities enable us to tackle a three-player problem using SDPs might be of independent interest.
Importantly, we are able to also show that these results can be adapted to show bounds for the lossy version of $\mathrm{QPV}_{\mathrm{BB84}}^{f}$. We do this by defining a new relaxation of our earlier SDP, which holds for the unentangled case, and show that the numerical bounds for this new SDP can be used to reprove a key lemma in [BCS22] for the case of our protocol. For instance, if the measurement error is very low, our results show that the single-qubit protocol $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  remains secure against attackers that share an amount of entanglement that is sublinear in the amount of classical bits, even when only, e.g. $51\%$ of the photons arrive at the honest party. The security proof even goes through in the case that the transmission of quantum information is slow, as long as classical information can be transmitted fast.
@@
The main result of this paper can be summarized as follows (see Theorems 3.11 and 5.6 for a formal version):
@@
(Informal). The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol is secure against attackers who pre-share a linear amount (in the size of the classical information) of qubits even if photon loss is almost 50% ($\eta>1/2$). Moreover, the family of its extensions $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ achieves security for smaller constant fraction of transmission rate (for $\eta>\frac{1}{m}$).
@@
Our work is directly related to the questions asked by Qi and Siopsis [QS15]. The most important difference between that work and ours, is that we are considering a more general security model. The attack analysis in [QS15] is based upon an assumption that attackers would immediately measure an incoming qubit using a projective measurement, and then communicate classically. By reducing to (extensions of) monogamy-of-entanglement games, we instead capture any quantum action of the attackers, including a round of quantum communication—it’s shown in [ABSVL22] that this stronger security model can in fact make a difference in the security. Critically, this also means that our method of analysis extends to the case where a classical function is combined with a single qubit, and our results make it possible to also prove entanglement bounds, instead of just bounds against unentangled protocols. Finally, our results can be reapplied in other settings that can be described by a monogamy-of-entanglement game.
There are two other recent papers by Allerstorfer, Buhrman, Speelman, and Verduyn Lunel [ABSL22, ABSVL22] that study the role of loss in quantum position verification. Those works provide many new results on the capabilities and limitations of fully loss-tolerant protocols.
Our work builds upon the (unpublished) results of Buhrman, Schaffner, Speelman, and Zbinden [Spe16b, Chapter 5]. That work took the initial step to bound QPV protocols in the lossy case using SDPs. However, its analysis was incomplete and numerically significantly less tight, e.g. its bounds on $\mathrm{QPV_{BB84}}$ were a factor two worse than the current work. Additionally, we significantly extend its applicability by extending the technique to also show entanglement bounds for the $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  protocol [BFSS13, BCS22].
2 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol
@@
Let $\eta$ denote the transmission rate of the qubits sent from $V_{0}$ to P. We define one round of the lossy-BB84 QPV protocol, denoted by $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, as follows:
@@
    $V_{0}$ and $V_{1}$ secretly agree on random bits $v,x\in\{0,1\}$. Then, $V_{0}$ prepares the qubit state $\left|\phi\right\rangle=H^{x}\left|v\right\rangle\in\{\left|0\right\rangle,\left|1\right\rangle,\left|+\right\rangle\,\left|-\right\rangle\}$, where $H$ denotes the Hadamard transformation.
@@
    $V_{0}$ and $V_{1}$ send $\left|\phi\right\rangle$ and $x$ to P, respectively, with both signals propagating at the speed of light in vacuum. The two verifiers coordinate so that the qubit and the classical bit arrive to P simultaneously.
@@
    Immediately, P measures the received qubit in the basis $x$ and broadcasts her outcome, either 0 or 1, to $V_{0}$ and $V_{1}$. If she did not receive the qubit, i.e.  the photon was lost, she sends $\perp$. Therefore, the possible answers from P are $v_{\textsf{P}}\in\{0,1,\perp\}$.
@@
      $V_{0}$ and $V_{1}$ receive their respective answers at the time corresponding with the claimed location, and they are equal, i.e.  both receive the same $v_{\textsf{P}}$, then, if
      • $v_{\textsf{P}}=v$, the verifiers record ‘Correct’, denoted by ‘c’,
      • $v_{\textsf{P}}=1-v$, the verifiers record ‘Wrong’, denoted by ‘w’,
      • $v_{\textsf{P}}=\perp$, the verifiers record ‘No photon’, denoted by $\perp$’,
@@
See Fig. 1 for a schematic of the original $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol. By setting $\eta=1$ and $p_{\mathrm{err}}=0$, one recovers the original $\mathrm{QPV}_{\mathrm{BB84}}$ protocol.
[…]
@@
Figure 1: Schematic representation of steps 2 and 3 of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol, where straight lines represent classical information and undulated lines represent quantum information.
In order to accept or reject the location, the verifiers run $r$ sequential rounds of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$. Let $r_{\textsc{C}}$, $r_{\textsc{W}}$, $r_{\perp}$ and $r_{\lightning}$ denote the number of times that the verifiers output ‘Correct’, ‘Wrong’, ‘No photon’ and ‘Abort’, respectively, after the $r$ rounds. Given transmission rate $\eta$ and error probability $p_{err}$, the verifiers expect
$\begin{split}r_{\textsc{C}}&\approx r\eta(1-p_{err}),\hskip 42.67912ptr_{\textsc{W}}\approx r\eta p_{err},\\ r_{\perp}&\approx r(1-\eta),\hskip 62.59596ptr_{\lightning}=0.\end{split}$ (1)
The symbols $\approx$ are due to the fact that in an implementation, the above values would only be reached for $r\to\infty$. To make a decision based on the observed outcomes, one must first define what it means for them to be sufficiently “close” to the expected values in (1). To this end, we introduce a binary test (Definition 2.2) that determines whether the prover’s location should be accepted or rejected based on the received data.
[…]
$\Gamma_{r}=\sum_{i=1}^{r}T_{i}(\textbf{a}_{i}),$ (2)
be the total score after $r$ rounds. Fix a parameter $\varepsilon_{\textrm{h}}>0$, which determines the confidence level of the test that we introduce. That is, it will succeed with probability at least $1-\varepsilon_{\textrm{h}}$ when interacting with an honest prover.
Definition 2.2.
Let $\varepsilon_{\textrm{h}}>0$. For the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol executed sequentially $r$ times, we define the acceptance test $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$, also referred to as the decision criterion, as follows: the verifiers accept the prover’s location if
$\Gamma_{r}\geq r(\alpha-\delta),$ (3)
where $\delta=\sqrt{\frac{\cos^{4}\frac{\pi}{8}\ln(1/\varepsilon_{\textrm{h}})}{r}}$. Otherwise, they reject.
We will next see that the test $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ is complete: an honest implementation of the protocol will be accepted, except with negligible probability. For an honest prover ($hp$), we have that, for every round $i$,
and therefore, $\mathbb{E}[\Gamma_{r}^{hp}]=r\alpha$. For simplicity, we will assume the dependence on $\eta$ and $p_{err}$ in $\alpha$ implicit. Then, by Hoeffding’s inequality [Hoe63], see Lemma 2.3, an honest prover will be accepted except with small probability at most $\varepsilon_{\textrm{h}}$.
Lemma 2.3.
Let $T_{1},\ldots,T_{r}$ be independent bounded random variables with $T_{i}\in[x_{a},x_{b}]$, for all $i\in\{1,\ldots,r\}$, with $-\infty<x_{a}\leq x_{b}<\infty$. Then, for all $\delta\geq 0$, the following holds:
$\mathrm{Pr}\left[\frac{1}{r}\sum_{i=1}^{r}(T_{i}-\mathbb{E}[T_{i}])\geq\delta\right]\leq e^{-\frac{2r\delta^{2}}{(x_{b}-x_{a})^{2}}}.$ (5)
In addition, we will see that for any attackers in the No-PE model—which imposes only the restriction that no entanglement is shared prior to the execution of the protocol in each round—the test $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ is sound: any attackers in this model will be rejected with high probability—in particular, we will show exponentially high probability (in $r$).
Remark 2.4.
This test is engineered from the analysis of the correlations attainable by attackers in Section 2, ensuring that they are rejected except with negligible probability; see the proof of Theorem 2.11 for details of its construction.
We say that a protocol is secure in a given model if it admits a test that is both complete and sound (for any adversaries acting according to the model). Next, we will show that $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ fulfills both conditions for a certain range of transmission rate $\eta$ and qubit error $p_{err}$ and thus showing that $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ is secure.
In order to prove security for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, we will first show that, in a round, for a range of values of $\eta$ and $p_{\mathrm{err}}$, any adversaries in the No Pre-shared Entanglement (No-PE) model are unable to reproduce the outcome probabilities specified in (1). This provides intuitive evidence that attackers cannot mimic the behavior of an honest prover, and that their actions can be statistically distinguished. We will later show that the test $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ achieves this.
For the security analysis, we will consider the purified version of $\mathrm{QPV}_{\mathrm{BB84}}$, which is equivalent to it. In this version, instead of $V_{0}$ sending BB84 states, $V_{0}$ prepares an EPR pair $\left|\Phi^{+}\right\rangle_{VP}$ and sends the register $P$ to the prover and keeps register $V$. At a later moment, $V_{0}$ performs the measurement $\{H^{x}\left|v\rangle\langle v\right|_{V}H^{x}\}_{v\in\{0,1\}}$ in his register $V$. In this way, the verifiers delay the choice of basis in which the qubit is encoded, which, in contrast to the above prepare-and-measure version, will make any attack independent of the state sent by $V_{0}$.
@@
Whereas it is well-known that adversaries sharing an unbounded amount of entanglement can always successfully break any QPV protocol [BCF+14], the proof of the security of the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol under attackers that do not pre-share entanglement [BCF+14] opened a branch of study, motivated by adversarial models that restrict attackers in a more realistic way. The most generic attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ (purified version) in the No-PE model consists of:
@@
    Alices and Bob prepare an arbitrary quantum state $\sigma_{A_{0}}$, and Alice holds it. Since the attackers do not pre-share entanglement, any quantum operation that Bob could later perform as a function of $x$ can be included in Alice’s operation (see e.g. [BCF+14, TFKW13]).